HUBBLE SPACE TELESCOPE DAILY REPORT #2845 PERIOD COVERED: 0000Z (UTC) 04/05/01 - 0000Z (UTC) 04/06/01 Daily Status Report as of 096/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2/STIS/CCD 8679 (Constraining the Age of the Oldest Stars from the White Dwarf Cooling Sequence in M4) The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to search for the oldest and hence coolest white dwarfs {WDs} in M4, the nearest Galactic globular cluster to the Sun. New models for cooling hydrogen white dwarfs predict that their colors should become bluer, rather than redder, with increasing age. The goal is to test these models and at the same time constrain the age of M4. The proposal completed with no problems. 1.2 Completed Two Sets of WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed nominally. 1.3 Completed STIS/CCD 8864 (CCD Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks for the CCD. The proposal completed nominally. 1.4 Completed Two Sets of WF/PC-2 8867 (Gamma-Ray Bursts: Discovering The Progenitors And Understanding The Explosion - Visits A0-R0) The WF/PC-2 was used to observe a gamma-ray burster, GRB010222. Gamma-ray burst astronomy, one of the most active and exciting frontiers in astrophysics, is now entering a critical stage -- with dramatic leaps in understanding of these events, as well as new discoveries. Improvements in triggering and positioning accuracy provided by the SAX and HETE-2 gamma-ray satellites will allow entirely new classes of events to be studied. Given the recent progress in this field, the proposers are now in a position to design precision, broadband measurements that can provide quantitative information on the as-yet unknown energy sources, the explosion geometry, and the surrounding medium. In particular, the growing evidence of an intimate connection between SNe and GRBs can be definitively tested. The proposal completed with no reported problems. 1.5 Completed Four Sets of STIS/CCD 9248 (Probing the Large Scale Structure: Cosmic Shear Observations) The Space Telescope Imaging Spectrograph (CCD) was used to measure the distortion of light bundles from distant galaxies that probe the statistical properties of the intervening inhomogeneous {dark} matter distribution. Its tidal gravitational field distorts the observable image shapes thereby causing a coherent ellipticity pattern {Cosmic Shear}. The observations completed nominally. 1.6 Completed WF/PC-2 8816 (Cycle 9 UV Earthflats) The WF/PC-2 was used to obtain sequences of Earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set and in order to monitor flat field stability. There were no reported problems. 1.7 Completed WF/PC-2 8720 (Masses and Multiplicity of Nearby Free- floating Methane and L Dwarfs) The WF/PC-2 was used to observe 50 very-low-mass objects in the solar neighborhood with spectral types of L0 and later {including several dwarfs with Methane absorption bands in their atmospheres}. These objects will be observed in two filter bands with the aim to identify close companions, measure their colors, and to obtain first epoch data of the newly discovered binaries. The observations completed nominally. 1.8 Completed STIS/CCD 8851 (Sparse Field CTE test {Cycle 9}) The Space Telescope Imaging Spectrograph (CCD) was used to perform CTE measurements that are made using the "sparse field test", along both the serial and parallel axes. The proposal completed with no reported anomalies. 1.9 Completed STIS/CCD 8865 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 9 Successful: 9 Scheduled Re-acquisitions: 7 Successful: 7 2.2 FHST Updates: Scheduled: 18 Successful: 18 2.3 Operations Notes: The SSR EDAC error counter was reset three times, using ROP SR-1A. The STIS MCE-1 reset at 095/113938Z while the low voltage was on and while inside an SAA interval. The STIS flight software error counter was cleared per ROP NS-12. MAMA-2 was recovered at 095/2353Z via normal SMS commanding. The CCS engineering status buffer limits were adjusted at 096/0727Z, per ROP DF-1A. A TTR was written when a re-transmit was required at 096/0713Z during a 486 command load uplink. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW